40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1401     Entry time: Fri Mar 13 20:23:37 2009
Author: Yoichi 
Type: Update 
Category: LSC 
Subject: AO path transfer function with X-arm locked 
I measured the AO path transfer function while the X-arm is locked with the POX PDH signal.
The POX-I signal was already connected to the input 1 of the CM board. So I injected a signal from the EXC-B channel of the board and measured the transfer function from TP2B to TP1A. To open the loop, I disabled the switch befor the EXC-B.
The attached plot shows the measured transfer function.
There is a bump around 2kHz, which can also be seen in the AO path TF posted in elog:1399, but not the large structure at around 3.8kHz.
The 3.8kHz structure is probably created by the feedback.
Attachment 1: AOPath-Xarm.png  81 kB  | Show | Show all
ELOG V3.1.3-